an object moving at 15m/s slows uniformly at the rate of 3m/s each second for a time of 2 sec..determine final velocity average velocity and distance moved


1
Expert's answer
2021-05-09T15:08:22-0400

Gives

"u=15m\/sec"

acceleration (a)="-\frac{3}{1} m/sec^2"

acceleration (a)=-3"m\/sec^2"

Time(t)=2sec

Final velocity

"v=u+at"

"v=15-3\times2"


"v=9m\/sec"

Distance move

"s=ut+\frac{1}{2}at^2"

"s=15\times2+\frac{1}{2}(-3)\times2^2"

s=24 m


average velocity

Vave="\frac{displacement}{time}"

vave="\frac{24-0}{2-0}=12m/sec"